### v2

# Universal Phase Transitions of Matter in Optically Driven Cavities

Abstract
Optical cavities have been widely applied to manipulate the properties of solid state materials inside it. We propose that in systems embedded within optical cavities driven by incident pump light, the pump induces generic phase transitions into new non-equilibrium steady states. This effect arises from the ponderomotive potential, the effective static potential exerted by the pump on the low energy degrees of freedom, which exhibits a universal step-like structure that pushes the matter degrees of freedom in the direction that red-shifts the cavity photon modes. For a two dimensional electron liquid in a driven cavity, this step-like potential pushes the electron density to jump to a smaller value so that a hybrid cavity photon mode is red shifted to slightly below the pump frequency. Similarly, for a dirty superconductor in such a driven cavity, this potential acts on the superconducting order parameter and causes a first order phase transition to a new steady state with a smaller gap. By realistic electromagnetic modeling of the cavity that includes all cavity modes, we construct the non-equilibrium phase diagrams for experimentally relevant devices.
An emerging platform for tuning the properties of solid state materials is to couple them to optical cavities [1, 2, 3, 4, 5, 6, 7], which enables precise control of light-matter interactions in confined geometries. Apart from the hybridization between collective modes and cavity photons into polaritons [1, 4, 8, 9], researchers are actively exploring the possibility that the cavity photon fluctuations may even manipulate phases of matter such as superconductivity [10, 11, 12, 13, 14, 15], ferroelectricity [16, 17, 18], excitonic order [19, 20], magnetism [21, 22], metal-insulator transitions [23], superradiance [24], and topology [25, 26, 27, 28, 29, 30]. Unfortunately, because of the smallness of the quantum and thermal fluctuations of cavity photons, it is experimentally challenging to realize cavity engineering of equilibrium many-body phases [25, 23, 31, 28, 29, 32]. On the other hand, one may greatly enhance the cavity photon fields by pumping it with incident light within optically driven cavities, so that the material inside it could be much more effectively controlled in the non-equilibrium regime [33, 34, 35, 36, 37, 38, 39, 40, 41, 42].
In this letter, we investigate the non-equilibrium steady states of many-body systems (the ‘sample’ in Fig. 1(a)) placed within optically driven Fabry–Pérot (F-P) cavities. By an electromagnetic approach that analytically accounts for the coupling of matter to all cavity photon modes without artificial parameters, we uncover a universal step-like ponderomotive potential [43] on the sample generated by the pump, leading to generic phase transitions for almost all systems. Non-equilibrium phase diagrams are predicted for the two dimensional electron gas (2DEG) and dirty superconductors to be verified by experiments.
Figure 1: (a) Schematic of a Fabry–Pérot cavity of thickness $h$ with the sample (an ultra-thin film or a strictly two-dimensional material) inside it. The mirror has the reflection and transmission coefficients $R(\omega)$ and $T(\omega)$. The red color scale represents the in-plane electric field of an anti-node cavity photon mode. (b) The dressed cavity photon frequencies $\omega_{\text{N}}$ at zero in-plane momentum as functions of $\phi$, the slow degree of freedom on the sample. The red color scale represents the ponderomotive force on $\phi$, which peaks at $\phi_{\text{c}}$ when the cavity photon is in resonance with the pump $\omega$. (c) The ponderomotive potential $V_{\text{P}}$ and ponderomotive force $F_{\text{P}}$ felt by $\phi$. (d) The typical total potential felt by $\phi$ for $E_{\text{p}}=0$ (black dashed line), $E_{\text{p}}=E_{\text{c}}$ (red dashed line), and $E_{\text{p}}>E_{\text{c}}$ (solid red line).
We consider a symmetric F-P cavity driven by the uniform pump light $\mathbf{E}_{\text{p}}e^{-i\omega t}+c.c.=-\partial_{t}\mathbf{A}_{\text{p}}$ represented by the electromagnetic vector potential $\mathbf{A}_{\text{p}}=\mathbf{A}_{\text{p0}}e^{-i\omega t}+c.c.$ with the frequency $\omega$, as shown in Fig. 1(a). The sample inside it is modeled as a two dimensional (2D) plane representing a 2D thin film or an atomic layer. The generic Lagrangian for such a system could be written as
$L=L_{\text{M}}[X,\phi,\mathbf{A}]+L_{\mathrm{EM}}[\mathbf{A},\mathbf{A}_{\text{p}}]\,.$ (1)
Here $L_{\text{M}}$ is the Lagrangian for the matter degrees of freedom in the sample: the fast field $X$ and the slow field $\mathbf{\phi}$ [43] which are coupled to the vector potential $\mathbf{A}(\mathbf{r},t)$, the photonic degree of freedom inside the cavity. The slow field $\phi$ represents the degrees of freedom whose dynamics is much slower than the driving frequency [43], e.g., a soft lattice distortion in an insulator, or the slow component of the order parameter in a superconductor. To grasp a physical picture, we encourage the readers to think of $\phi$ as the electronic density $n$ of a metal in the following. The fast field $X$ are the degrees of freedom that oscillate at the driving frequency or its higher harmonics, such as the high energy electronic transitions.
The term $L_{\mathrm{EM}}=L_{\mathrm{cavity}}+L_{\mathrm{pump}}$ contains the quadratic Lagrangian $L_{\mathrm{cavity}}$ for the cavity photon modes and $L_{\mathrm{pump}}$ for their linear coupling to the external pump field $\mathbf{A}_{\text{p}}$. They lead to the Maxwell equations and the mirror’s transmission/reflection coefficient $R(\omega)$$T(\omega)$ that encode the information of all the cavity photons and their coupling to the pump. It is convenient to define the dimensionless coefficients:
$\alpha^{\prime}$ $=\frac{\alpha}{1-i\beta\sigma/c},\quad\alpha=\frac{Te^{i\frac{kh}{2}}}{1-Re^{ikh}}$ (2)
where $h$ is the cavity thickness, $\alpha^{\prime}(\omega)$ is the effective linear transmission coefficient defined as the ratio between the electric field $E$ on the sample plane and the externally incident electric field $E_{\text{p}}$, while $\alpha(\omega)$ denotes the bare transmission coefficient in absence of the sample, see Supplemental Material (SM) Sec. I [44]. Here we have restricted to the case that the sample is in the middle of the symmetric cavity and the pump light is normally incident with the wave vector $k=\omega/c$. The dimensionless factor $\beta(\omega)$ has the meaning of the radiation coefficient inside the cavity viewed from the sample. It approaches $\beta\approx 2\pi\tan(kh/2)$ for a perfect cavity with $R\rightarrow-1$. Therefore, both $\alpha$ and $\beta$ contain resonant poles from bare anti-node cavity photons at $kh=(2N+1)\pi$. Through its 2D optical conductivity $\sigma(\omega)$ in Eq. (2), the sample hybridizes with bare cavity photons and push their eigen frequencies to those of the physical cavity photons labeled by $\omega_{\text{N}}$, which will be referred to as ‘cavity photons’ in the following. They appear as the poles of $\alpha^{\prime}$ determined by the zeros of its denominator (only anti-node cavity photons are involved).
Step-like ponderomotive potential—The coherent pump field drives the cavity photon which in turn drives the sample, and generates an effective static force on the slow degrees of freedom $\phi$, the ponderomotive force $F_{\text{P}}(\phi)$ [43, 46, 47, 48, 49, 50, 51, 52, 53, 54, 46, 55, 56, 57, 58, 59, 60, 61, 62]. Formally, the total force on $\phi$ is $F=-\langle\partial_{\phi}L\rangle$ where the bracket denotes time average for classical systems, and path-integral average over all paths of the fast fields $X$ and $A$ in the quantum mechanical case. By integrating out the fast degrees of freedom $X$ and $A$ using the Keldysh path integral [63, 64, 65, 66], one may obtain the low energy effective field theory $L[\phi]=L_{0}[\phi]+V_{\text{P}}[\phi]$ for $\phi$, whose potential term reads [43]
$V[\phi]=V_{0}[\phi]+V_{\text{P}}[\phi],\qquad F_{\text{P}}=-\partial_{\phi}V_{\text{P}}\,.$ (3)
Here $V_{0}[\phi]$ is the equilibrium term that exists even at zero pump field and $V_{\text{P}}[\phi]$ is the ponderomotive potential induced by the pump field $A_{\text{p}}$. At the mean field level, the configuration of $\phi$ that minimizes the potential landscape $V[\phi]$ gives the non-equilibrium steady state.
In the steady state, the cavity photon field must be oscillating periodically in time with a classically coherent value, the ‘mean field’, together with small quantum and thermal fluctuations. Within linear response, the classical value is simply $Ae^{-i\omega t}+c.c.$ where $A=\alpha^{\prime}(\omega)A_{\text{p}}$ is found from Maxwell’s equations using the optical conductivity $\sigma(\omega,\phi)$ at a fixed $\phi$. Therefore, the sample is now periodically driven by the cavity field $A$. The force $F=-\langle\partial_{\phi}L\rangle=-\langle\partial_{\phi}L_{\text{M}}\rangle_{X}$ on $\phi$ could be computed from $L_{\text{M}}$ in Eq. (1) with the classical cavity field $A$ treated as the external driving field, which yields the ponderomotive force $F_{\text{P}}$. The effective potential generated by the vacuum fluctuations of $A$ is a much weaker effect and is incorporated into the $V_{0}[\phi]$ part of Eq. (3), see SM Sec. IIC for an estimation [44]. In the absence of dissipation, meaning that there is no energy absorption within the sample, the driven sample belongs to case 1 below Eq. 4 of Ref. [43], and the ponderomotive force must be equal to that predicted by Eq. 4 there. At second order in the pump field, it is simply related to the equilibrium optical conductivity $\sigma(\phi,\omega)$ of the sample as
$F_{\text{P}}=\frac{i\omega}{c^{2}}|A|^{2}\partial_{\phi}\sigma=\frac{i\omega}{c^{2}}|\alpha^{\prime}|^{2}A_{\text{p}}^{2}\partial_{\phi}\sigma$ (4)
where $\omega>0$, $\alpha^{\prime}$ is from Eq. (2) and $\sigma$ is purely imaginary in the absence of dissipation, see Appendix B for detailed derivation.
One is now ready to observe that the ponderomotive potential from Eq. (4) has a universal step-like structure shown in Fig. 1(c). As $\phi$ varies, the optical conductivity of the sample changes, and thus the cavity photon frequencies $\omega_{\text{N}}(\phi)$ are shifted according to the poles of $\alpha^{\prime}$, as shown schematically in Fig. 1(b). Suppose that a cavity photon $\omega_{2}(\phi)$ is closely above the pump frequency $\omega$ when $\phi=\phi_{0}$ and red shifts with decreasing $\phi$ as in Fig. 1(b), it may hit a resonance with the pump at a ‘resonance’ value of $\phi=\phi_{\text{c}}$. There the ponderomotive force and potential from Eq. (4) are dominated by the resonant pole of $|\alpha^{\prime}|^{2}$:
$F_{\text{P}}\approx-\frac{V_{\text{u}}}{\pi}\frac{\gamma_{\phi}}{(\phi-\phi_{\text{c}})^{2}+\gamma_{\phi}^{2}},\quad V_{\text{P}}\approx\frac{V_{\text{u}}}{\pi}\arctan\frac{\phi-\phi_{\text{c}}}{\gamma_{\phi}}$ (5)
where $\gamma_{\phi}$ is the width of the resonance in the $\phi$ direction [67], see SM Sec. II [44]. Therefore, the ponderomotive force has a Lorentzian peak at $\phi_{c}$ because the cavity photon field is resonantly enhanced here, and the ponderomotive potential has a step-like ‘arctan’ structure, as shown in Fig. 1(c). Amazingly, if the sample has no dissipation so that the only contribution to the linewidth of the cavity photon is its radiative loss $\gamma_{\text{r}}=\omega|T|^{2}$ due to a nonzero transmission coefficient of the mirror, the height of the potential jump is a universal value
$V_{\text{u}}=E_{\text{p}}^{2}\lambda/(4\pi)$ (6)
that depends only on the electrical field and vacuum wavelength $\lambda=2\pi/k$ of the incident light, see SM Sec. II [44]. It has the physical meaning of the 2D EM field energy density of the pump laser in a 2D layer of thickness equal to its wavelength. We note that this is a special mechanism brought by a cavity, which is absent in conventional light-induced phase transitions in free-space [68, 69, 70, 71].
The sign of the ponderomotive force means that it always pushes $\phi$ in the direction so that $\mathrm{Im}[\sigma]$ becomes lower, leading to red shifts of the cavity photons. Note that according to the pole of $\alpha^{\prime}$, if $\sigma$ itself contains no extra poles, the $N$th cavity photon mode frequency is constrained between $(2N)\pi c/h$ and $(2N+2)\pi c/h$. Therefore, at most one cavity photon could be possibly tuned by $\phi$ to cross the resonance with the pump, meaning that there is at most one ponderomotive step on the potential landscape $V(\phi)$.
Together with the equilibrium potential whose structure close to the equilibrium value $\phi_{0}$ is $V_{0}\sim(\phi-\phi_{0})^{2}$, the total potential in Eq. (3) looks like that in Fig. 1(d). Because of the step-like ponderomotive potential, a dip appears closely to the left of $\phi_{\text{c}}$, creating a new potential minimum. When the pump strength of $E_{\text{p}}$ exceeds a critical value $E_{\text{c}}$, the new minimum becomes lower than the original one so that a first-order phase transition from $\phi=\phi_{0}$ to $\phi=\phi_{\text{c}}$ occurs. Note that the cavity photon degree of freedom plays an essential role in inducing this phase transition by endowing the ponderomotive potential with the step-structure in Eq. (5).
In the following, we apply this theory to two systems in driven cavities, and demonstrate the first order phase transitions induced by the ponderomotive step potential in Eq. (5).
A 2DEG in a driven cavity [72] is the first example, as shown in Fig. 2(a). The 2DEG is electrically connected to a gate electrode outside of the cavity, so that its carrier density is allowed to change. It is described by the Lagrangian in Eq. (1) with the matter part being
$L_{\mathrm{M}}$ $=\int d^{2}r\bar{\psi}\left[-i\partial_{t}+\xi(\hat{\mathbf{p}}+\mathbf{A})\right]\psi+V_{\text{i}}+V_{\text{ee}}$ (7)
where $\bar{\psi}$ and $\psi$ are the creation and annihilation fermion fields for electrons containing the fast degrees of freedom, while the slow degree of freedom $\phi$ considered here is the uniform electron density $n=\langle\bar{\psi}\psi\rangle$ on the sample, and $\hat{\mathbf{p}}=-i\nabla$ acts on the fermion field. The $V_{\text{i}}$ term represents electron-impurity scattering, and $V_{\text{ee}}$ is the electron-electron interaction including with those in the gate. For notational simplicity, the Planck constant $\hbar$, the elementary charge $e$ and the speed of light $c$ are set to be 1 but will be restored when necessary.
Figure 2: (a) Schematic of the 2D electron gas in the driven cavity. The 2DEG is electrically connected to a gate outside of the cavity. (b) Phase diagram of the device on the plane of pump frequency $\omega$ and electric field $E_{\text{P}}$, where the color scale corresponds to the calculated $\Delta n$, showing a clear phase boundary. The red dashed line shows the analytical prediction for the phase boundary from Eq. (10) with the damping correction. The parameters are $m=0.01\,\mathrm{m_{e}}$, $n_{0}=3\times 10^{14}\,\mathrm{\mathrm{cm}^{-2}}$, $h=1.2\,\mathrm{\mu m}$, $C=0.1\,\mathrm{\mu\mathrm{F}/\mathrm{cm}^{-2}}$, $T^{2}=1-R^{2}=2\times 10^{-4}$, and $\gamma=1\,\mathrm{meV}$ for the Drude optical conductivity. (c) The blue curve is $n_{c}$ as a function of $\omega$. The black dashed line represents $n_{0}$ and the red zone highlights its difference from the possible new phase.
For a parabolic band $\xi(\mathbf{p})=p^{2}/(2m)-\mu$ without impurity scattering, there is a quick derivation of the ponderomotive force because the paramagnetic coupling between electrons and light is inconsequential, and the only relevant coupling is the diamagnetic term: $nA^{2}/(2m)$. Taking a derivative with respect to $n$, the ponderomotive force for the electron density is obtained as
$F_{\text{P}}$ $=-\frac{e^{2}}{2m}\langle A(t)^{2}\rangle=-\frac{e^{2}}{m}|\alpha^{\prime}|^{2}A_{\text{p}}^{2}$ (8)
in agreement with Eq. (4), where $\alpha^{\prime}$ is from Eq. (2). Indeed, this is what we have proved in Eq. (4) considering that the optical conductivity $\sigma=ine^{2}/(m\omega)$ is proportional to $n$. The electron-electron interaction does not modify this result owing to the Galilean invariance of a parabolic-band electron liquid in the clean limit [73, 74, 75, 76]. This ponderomotive force pushes the carrier density to smaller values so that the cavity photon frequency shifts down. Therefore, the ponderomotive potential felt by $n$ is simply the step-like potential in Eq. (5) and Fig. 1(c) with $\phi$, $\phi_{\text{c}}$ and $\gamma_{\phi}$ replaced by $n$, $n_{\text{c}}$ and $\gamma_{\text{n}}$ whose values are found from the poles of $\alpha^{\prime}$ in Eq. (2). The total potential is Eq. (3) with the equilibrium term being $V_{0}(n)=e^{2}(n-n_{0})^{2}/(2C)$ where $n_{0}$ is the equilibrium carrier density, and the charge stiffness $1/C=1/C_{0}+1/\nu$ is set by the capacitance $C_{0}=1/(4\pi d)$ between the sample and the gate and the electronic density of states $\nu$ near the fermi surface. Therefore, the potential landscape is Fig. 1(d) with $\phi$ being the carrier density $n$.
When the carrier density $n$ is shifted to the ‘resonant density’ set by the poles of $\alpha^{\prime}$ in Eq. (2):
$n_{\text{c}}=-\frac{c\omega m}{e^{2}\beta}\approx-\frac{c\omega m}{2\pi e^{2}\tan(kh/2)}\,$, (9)
a resonance between a cavity photon mode and the pump occurs and $A$ and $F_{\text{P}}$ reach their maxima as shown in Fig. 1(c). The resonant density $n_{c}$ from Eq. (9) is plotted as a function of pump frequency in Fig. 2(c). Note that $n_{\text{c}}(\omega)$ has periodic ‘divergences’ around $kh=2N\pi$ for $N=1,2,3,$..., and periodic zeros at $kh=(2N-1)\pi$ when the pump is at resonances with bare anti-node cavity photon modes. From Fig. 1(d), it is obvious that if $0<n_{\text{c}}<n_{0}$ (red region in Fig. 2(c)), the step-like drop of potential creates a new potential minimum at $n_{\text{s}}\lessapprox n_{\text{c}}$. Physically, $n_{\text{c}}<n_{0}$ means that there is a cavity photon mode moderately above the pump frequency, so that if the carrier density is pushed down, the photon red shifts and potentially hits a resonance with the pump, see Fig. 1(b).
Due to the sharpness of the step potential, the energy difference between the two minima is simplified to $V(n_{\text{s}})-V(n_{0})\approx V_{0}(n_{\text{c}})-V_{\text{u}}$. If the external pump field is stronger than the critical value $E_{\text{c}}$ set by $V_{\text{u}}(E_{\text{c}})=V_{0}(n_{\text{c}})$, i.e.,
$E_{\text{c}}$ $\approx e\sqrt{\frac{k}{C}}(n_{0}-n_{\text{c}})$ (10)
the energy of the new minimum becomes lower than the original one, and the system undergoes a first order phase transition to this state of lower carrier density. This result may be relevant to the transitions observed numerically in driven electron-phonon systems [77, 78]. Note that if $n_{\text{c}}>n_{0}$, the drop of ponderomotive potential occurs to the right of $n_{0}$ in Fig. 1(d), and would not create a new minimum.
A numerically exact phase diagram computed from minimizing Eq. (3) is shown in Fig. 2(b) where the color scale represents the (negative) change of carrier density $\Delta n=n_{0}-n_{\text{s}}$ of the driven state compared to the equilibrium state. As the pump field $E_{\text{p}}$ exceeds the phase boundary following the vertical arrow, the system undergoes a first-order phase transition to the new phase where the electron density shifts abruptly from $n_{0}$ to $n_{\text{s}}$ so that its relevant photon mode becomes near resonance with (but slightly lower than) the pump. If one follows the horizontal arrow in Fig. 2(b), no phase transition occurs. At the initial point where $n_{\text{c}}>n_{0}$, there is only one minimum of $V(n)$ located at $n_{\text{s}}\lessapprox n_{0}$. As the pump frequency decreases so that $n_{\text{c}}$ decreases to below $n_{0}$, the global minimum $n_{\text{s}}$ sticks to the left of $n_{\text{c}}$ and moves smoothly with it, see Fig. 1(d). This type of phase diagram resembles that of the liquid-gas transition whose phase boundary terminates at a critical point.
Note that for pumping frequencies closely below the equilibrium cavity photon modes labeled by $\omega_{\text{N}}$ in Fig. 2(c), the critical pump field $E_{\text{c}}$ actually approaches zero because it only takes a small amount of carrier density shift for the photon to red shift to resonate with the pump. Therefore, these are the best frequencies to achieve this new phase experimentally. To avoid over heating in an actual experiment, an ultrafast laser pulse may be applied with its parameters tuned in time following the horizontal and vertical arrows in Fig. 2(b). The former steers the system to the new minimum which manifests as a dip/peak of the reflected/transmitted photons, while the later does not because there is not enough time to overcome the energy barrier.
When there is a nonzero scattering rate $\gamma$ so that the optical conductivity becomes the Drude form $\sigma=ine^{2}/[m(\omega+i\gamma)]$, it turns out that Eqs. (8)(5) still hold, while the potential drop $V_{\text{u}}$ in Eq. (6) is suppressed by a factor of about $\gamma_{\text{r}}/(\gamma_{\text{r}}-2\gamma\sin kh)$ where $\sin kh<0$ in the phase transition region (see SM Sec. III [44]). The critical field is thus larger than Eq. (10) by the inverse square root of this factor.
Dirty superconductors—A more intriguing example is a BCS-type dirty superconductor, either a thin film or a 2D superconductor, in a driven THz cavity [72, 79, 80] with a configuration shown in Fig. 1(a). In this device, the sample is not electrically connected to a gate so that its charge density is not allowed to change. Instead, the order parameter $\Delta$ for the superconducting state can vary, which is the low energy degree of freedom of the system. The matter part of the Lagrangian density in Eq. (1) could be written as [81]
$\mathcal{L}_{\text{M}}=\bar{\Psi}$ $\begin{pmatrix}-i\partial_{t}+\xi(\hat{\mathbf{p}}+\mathbf{A})&\Delta\\ \Delta^{\ast}&-i\partial_{t}-\xi(-\hat{\mathbf{p}}+\mathbf{A})\end{pmatrix}\Psi$
$+$ $V_{\mathrm{dis}}(\mathbf{r})n(\mathbf{r})+\frac{1}{g}|\Delta|^{2},$ (11)
where $\Psi=(\psi_{\uparrow},\bar{\psi}_{\downarrow})^{T}$ is the fermion field for spin up and down, $\xi(\mathbf{p})=\varepsilon(p)-\mu$ is the electronic kinetic energy same as Eq. (7), $V_{\mathrm{dis}}$ represents the disorder potential, and $g$ is the s-wave attractive interacting strength between electrons.
Figure 3: (a) The total potential felt by the order parameter $\Delta$ of a superconductor in the device in Fig. 1(a) for $E_{\text{p}}=0$ (black dashed line), $E_{\text{p}}\approx E_{\text{c}}$ (blue dashed line), and $E_{\text{p}}>E_{\text{c}}$ (red solid line). (b) Numerically exact phase diagram of the device where the color scale represents $\delta\Delta=\Delta_{0}-\Delta_{\text{s}}$, the (negative) change of gap relative to its equilibrium value. A clear phase boundary is marked by the red solid curve, while the red dashed line is the analytical approximation to it. The parameters are $m=0.3\,\mathrm{m_{e}}$, $\tau=0.03\,\mathrm{ps}$, $\Delta_{0}=12\,\mathrm{K}$, $g\nu=0.5$, $n_{0}=1\times 10^{13}\mathrm{cm}^{-2}$, $h=5\,\mathrm{mm}$ and $R=0.99$.
We now study the ponderomotive force exerted by the pump $A_{\text{p}}$ on the order parameter $\Delta$ which is taken to be real without loss of generality. To capture the essential physics without introducing much complexity, we discuss the case of zero temperature with the equilibrium gap $\Delta_{0}$ so that there are no thermally excited quasi-particles. For a dirty superconductor, the Mattis-Bardeen theory [82] renders a good approximation to the optical conductivity, which at low frequencies ($\omega\ll 2\Delta$) is just the superfluid response
$\sigma(\omega)\approx i\frac{ne^{2}}{m}\frac{\pi\tau\Delta}{\omega}\,.$ (12)
Here $n$ is the total carrier density, $\tau$ is the disorder limited electronic mean free time, and the superfluid density $n\pi\tau\Delta$ is proportional to the gap. Therefore, similar to the case of a 2DEG, the oscillating electric field tends to push the gap to smaller values to reduce the superfluid density, so that the cavity photon gets red-shifted.
Formally, for a sub-gap pump $\omega<2\Delta_{0}$, there is no dissipation from quasiparticles or pair-breaking excitation. Therefore, Eq. (4) and Eq. (5) apply here. The ponderomotive force felt by $\Delta$ is simply $F_{\text{P}}=-\pi n\tau\frac{e^{2}}{m}A^{2}=-\pi n\tau\frac{e^{2}}{m}|\alpha^{\prime}|^{2}A_{\text{p}}^{2}$, and the ponderomotive potential $V_{\text{P}}$ is the step-like potential in Eq. (5) with $\phi$, $\phi_{c}$ and $\gamma_{\phi}$ replaced by $\Delta$, $\Delta_{\text{c}}$ and $\gamma_{\Delta}$ whose values are found from the poles of $\alpha^{\prime}$. The ‘resonance gap’
$\Delta_{\text{c}}=-\frac{1}{\pi n\tau}\frac{c\omega m}{e^{2}\beta}$ (13)
simply has an extra constant factor compared to $n_{\text{c}}$ in Eq. (9) so that its frequency dependence looks like Fig. 2(c). When the order parameter is shifted to $\Delta_{\text{c}}$, the cavity photon frequency is shifted to a point in resonance with the pump, and a step-like drop of $V_{\text{P}}$ occurs.
The equilibrium potential $V_{0}(\Delta)=\Delta^{2}/g-\nu\Delta^{2}\ln(\Lambda/\Delta)$ is from standard BCS theory [81, 83] with $\nu$ being the density of states and $\Lambda$ being the energy cut-off, whose minimum gives the equilibrium gap $\Delta_{0}=\Lambda e^{-1/(g\nu)-1/2}$. With the pump, the total potential $V(\Delta)=V_{0}+V_{\text{P}}$ is modified to that in Fig. 3(a). If $0<\Delta_{\text{c}}<\Delta_{0}$, the step-like drop of potential occurs to the left of $\Delta_{0}$ as shown in Fig. 3(a), creating a new energy minimum at $\Delta_{\text{s}}$. A strong enough pump would make the energy of the new minimum lower than the original one, inducing a first order phase transition of the superconducting order parameter from $\approx\Delta_{0}$ to $\Delta_{\text{s}}\lessapprox\Delta_{\text{c}}$. Similar to Fig. 2(b), the numerically exact phase diagram is shown in Fig. 3(b) which displays a clear phase boundary.
Since the superconductor introduces no extra loss to the cavity photon, the step-like drop of the ponderomotive potential is the same universal value $V_{\text{u}}=E_{\text{p}}^{2}/(2k)$. The critical field $E_{\text{c}}$ could be estimated simply by $V_{0}(\Delta_{\text{c}})-V_{0}(\Delta_{0})=V_{\text{u}}(E_{\text{c}})$ which gives the red dashed curve in Fig. 3(b), a decent approximation to the phase boundary. If $\Delta_{\text{c}}$ is close to $\Delta_{0}$ which happens for pump frequencies closely below the equilibrium cavity photons, the equilibrium potential could be approximated by a quadratic one so that the criterion becomes $\nu(\Delta_{\text{c}}-\Delta_{0})^{2}=V_{\text{u}}(E_{\text{c}})$ and renders a very small critical field $E_{\text{c}}=(\Delta_{0}-\Delta_{\text{c}})\sqrt{2k\nu}$ [84].
To conclude, we note that Eqs. (4) (5) for the dirty superconductor is only strictly correct at zero temperature so that there is no optical absorption from quasi-particles, while the Eliashberg effect [85, 86, 35] that enhances superconductivity dominates at temperatures close to the critical temperature, see Appendix C for nonzero temperature effects. Although we focused on Fabry–Pérot cavities for analytical clarity, the other types of cavities such as plasmonic/polaritonic nano cavities can also be treated within our framework with the coefficient $\alpha^{\prime}(\omega,\phi)$ replaced by the appropriate expressions. Future directions include exploring other new phases induced by this potential in driven cavities, the effect of fluctuations beyond the mean-field level, and the force induced by vacuum fluctuations in equilibrium cavities [25, 27, 28, 29, 30, 15].
Acknowledgements.
This work is supported by the National Key Research and Development Program of China (2022YFA1204700), the National Natural Science Foundation of China (Grants No. 12374291 and No. 12421004), Beijing Natural Science Foundation (Z240005), and the startup grant from Tsinghua University. We thank M. Yarmohammadi, Q. Xiong, T. Qin and T. Xiao for helpful discussions.
Data availability—Numerical codes and data for plots in this paper are available online [87].